How many joker cards are there in a deck of playing cards?
There are usually two joker cards in a deck of playing cards. They can used to replace a damaged, soiled or lost playing card or included in games like Old Maid.
What is the formula used to calculate the probability of arranging 1 to n in the correct order?
There are n! = 1*2*3*...*n possible orders and out of these, only one is correct. So the probability is 1/n!
There are n! = 1*2*3*...*n possible orders and out of these, only one is correct. So the probability is 1/n!
There are n! = 1*2*3*...*n possible orders and out of these, only one is correct. So the probability is 1/n!
There are n! = 1*2*3*...*n possible orders and out of these, only one is correct. So the probability is 1/n!
How do you obtain a probability distribution?
Find all the possible outcomes and the probabilities associated with each. That information comprises the probability distribution.
Find all the possible outcomes and the probabilities associated with each. That information comprises the probability distribution.
Find all the possible outcomes and the probabilities associated with each. That information comprises the probability distribution.
Find all the possible outcomes and the probabilities associated with each. That information comprises the probability distribution.
Why do you need probability in real life?
Because there are many events whose outcomes cannot be determined. However, using probability it may be possible to make a good estimate as to the outcome.
Because there are many events whose outcomes cannot be determined. However, using probability it may be possible to make a good estimate as to the outcome.
Because there are many events whose outcomes cannot be determined. However, using probability it may be possible to make a good estimate as to the outcome.
Because there are many events whose outcomes cannot be determined. However, using probability it may be possible to make a good estimate as to the outcome.
What is the probability for someone going to college?
The answer is pretty close to zero. This is because out of the millions of people across the world, a very small proportion make it to college.
Incidentally, "college" is not a term that is used consistently. For example, Eton (the famous English "public" school in Windsor, is actually called Eton College. On the other hand, the London School of Economics and Political Science is for graduate (and post-grad) studies. As a result, the question is somewhat open to interpretation.
The answer is pretty close to zero. This is because out of the millions of people across the world, a very small proportion make it to college.
Incidentally, "college" is not a term that is used consistently. For example, Eton (the famous English "public" school in Windsor, is actually called Eton College. On the other hand, the London School of Economics and Political Science is for graduate (and post-grad) studies. As a result, the question is somewhat open to interpretation.
The answer is pretty close to zero. This is because out of the millions of people across the world, a very small proportion make it to college.
Incidentally, "college" is not a term that is used consistently. For example, Eton (the famous English "public" school in Windsor, is actually called Eton College. On the other hand, the London School of Economics and Political Science is for graduate (and post-grad) studies. As a result, the question is somewhat open to interpretation.
The answer is pretty close to zero. This is because out of the millions of people across the world, a very small proportion make it to college.
Incidentally, "college" is not a term that is used consistently. For example, Eton (the famous English "public" school in Windsor, is actually called Eton College. On the other hand, the London School of Economics and Political Science is for graduate (and post-grad) studies. As a result, the question is somewhat open to interpretation.
What is probability of rolling a 2 or smaller?
Assume we want to determine the probability of rolling a 2 or smaller of a fair die (If it isn't, then some probabilities of getting a certain number must be different from the usual).
List the possibilities for such event. We have:
P(2 or smaller) = P(2) + P(1)
The probability of obtain a number on a die is 1/6. Therefore, the probability of obtaining a 2 or smaller on a fair die is 1/6 + 1/6 = 2/6 = 1/3.
What is the describing of the complementary event and find its probability?
Suppose there is an event A and the probability of A happening is Pr(A). Then the complementary event is that A does not happen or that "not-A" happens: this is often denoted by A'.
Then Pr(A') = 1 - Pr(A).
Suppose there is an event A and the probability of A happening is Pr(A). Then the complementary event is that A does not happen or that "not-A" happens: this is often denoted by A'.
Then Pr(A') = 1 - Pr(A).
Suppose there is an event A and the probability of A happening is Pr(A). Then the complementary event is that A does not happen or that "not-A" happens: this is often denoted by A'.
Then Pr(A') = 1 - Pr(A).
Suppose there is an event A and the probability of A happening is Pr(A). Then the complementary event is that A does not happen or that "not-A" happens: this is often denoted by A'.
Then Pr(A') = 1 - Pr(A).
How many different ways can you arrange the letters of the word chocolate?
chocolate = 9 letters, where o and c are repeated 2 times.
There are 9!/(2!2!) = 90,720 ways.
What is the probability that the spinner will land on the number 1?
It is the proportion of the spinner's perimeter that is occupied by the section (or sections) with a value of 1.
What is the probability of 1 out of three businesses succeed if it succeeds 60 percent of the time?
Pr(1 out of 3 succeed) = Pr(1 succeeds and 2 fail)
= 3*(0.6)*(0.4)2 = 0.288
There is no such term.
The regression (or correlation) coefficient changes as the sample size increases - towards its "true" value. There is no measure of association that is independent of sample size.
What is the probability of rolling 7 with 2 dice?
1:6. It's easy to see why if you think about it a certain way... First, roll one die. Whatever number you roll, there is exactly one number you can roll on the other die to total seven. The odds that you'll roll that one other number is 1:6 (because the die has 6 sides).
There are 52 cards in a deck.
4*4 = 16 of those are either a seven,jack, queen or king.
.
16 out of 52 = 16/52 = 0,307692307692307 = 30,77 %
What is the probability that a random couple with three children will have all girls?
Assuming that children of either gender are equally likely, the answer is (1/2)3 = 1/8
3:4 or 75%
What is the probability of drawing either a Jack and a 5 from a deck of cards?
If one card is drawn at random, the probability is 2/13.
27/64
There are four suits in a deck of cards with equal numbers of cards per suite, so the chance of drawing a heart is 1/4 and a spade is 1/4 too.
To draw a heart, replace it, then draw a spade, 1/4x1/4 = 1/16, or a 6.25% chance.
It is 1/169 or 0.0059, approx.
Which is not a possible vaule for a probability?
Probability value can only be real number between 0 to 1, therefore, it cannot be greater than 1 or less than 0 or a complex number.